An ideal gas undergoes a mechanically reversible process (constant pressure, constant temperature and adiabatic process. The gas entering a T1=650K and P1=10bar decreases its temperature at constant pressure where V2=2.91x10-³m³

Then, it went to isothermal process to decrease its pressure. Finally, the gas returns to its initial state. Take Cp=7/2 and Cv==5/2 R. 


Calculate the following:


a. T2,P3, V1 and V3


b. Q, W,ΔU and ΔH​


1
Expert's answer
2021-11-19T09:16:01-0500

T1=650K


P1= 10bar


V2=2.91x10-3


Constant P and T


V1=?


Pdv= nRTPdv=nRT


10(V2-V1) = 8.314×650


2.91x10-2 -10v1 =5404.1


V1=5.4×103